OpenStudy (anonymous):

Medals!!! There are 19 parts bran flakes and 1 part raisins in a certain cereal. If the volume of the box is 32 oz, how many ounces of each food are in the box?

OpenStudy (anonymous):

20oz=19oz bran + 1oz raisin

OpenStudy (anonymous):

1oz= 19/20 oz bran + 1/20 oz raisin

OpenStudy (anonymous):

32oz= (19/20)*32oz bran + (1/20)*32oz raisin

OpenStudy (anonymous):

What is the complete answer?

OpenStudy (anonymous):

im not allowed to put the final answer. Sorry. this program warned me once

OpenStudy (anonymous):

Oh. So what do I do next?

OpenStudy (anonymous):

I'm stuck here: 20x=32 oz.

OpenStudy (anonymous):

(19/20)*32oz =bran (1/20)*32oz =raisin

OpenStudy (anonymous):

Are 19/20 fractions?